OpenStudy (iamstark):

A hollow sphere has radius 6.4 m. Minimum velocity required by a motor cyclist at the bottom to complete the circle will be

OpenStudy (iamstark):

@iambatman @IrishBoy123

OpenStudy (michele_laino):

hint: when the motor cyclist is at the top of the circumference, then the subsequent condition must hold: weight= centrifugal force \[\huge mg = m\frac{{v_{\min }^2}}{R}\] where R is the radius of the circumference, and Vmin is the requested speed

OpenStudy (iamstark):

Boing!!!! That means the velocity should be 8m/s !! But the answer is 17.7 m/s ;-;

OpenStudy (iamstark):

Oh !!!!! I get it now, I forgot to consider the fact that the speed will keep decreasing as the biker climbs the sphere. So that means he must start with a speed greater than 8 m/s from bottom. I can use equations of kinematics for that \(\sf 8^2 = u^2 -2 \times -2\times g \times 2R\) !!! yaY !!